OpenStudy (anonymous):

What value of z makes the equation true? 2.5 = –6.19 + z

OpenStudy (anonymous):

@TheAsker2002

OpenStudy (anonymous):

@dat

OpenStudy (anonymous):

do opposite operation

OpenStudy (anonymous):

the answer is still 8.79

OpenStudy (anonymous):

wait nvm

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

did i type 8.69 on the last one? my finger must have slipped.

OpenStudy (anonymous):

no it is 8.69

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

hey! dood i know you are lying you look things up get the answers and get the reword now stop foiling around @basilalter

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

@mr_basketball27

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

what am I looking up? this is literally basic math

OpenStudy (anonymous):

this is probably the easiest question I've ever answered. ever.

OpenStudy (anonymous):

whatever get out of here cheater

OpenStudy (anonymous):

you can ask everyone on here and they will give you the same answer.

OpenStudy (anonymous):

what am I cheating on?

OpenStudy (anonymous):

\[2.5= -6.19+2\] \[+6.19\] \[=8.69\]

OpenStudy (anonymous):

@mr_basketball27 thanks

OpenStudy (anonymous):

rilly? ok thanks i have 8 more

OpenStudy (anonymous):

really

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

oint K is graphed on a number line at 3. Point L is 11 units away from point K on the number line. What could be the coordinates of point L? Choose all answers that are correct. A. –14 B. –8 C. 7 D. 14

OpenStudy (anonymous):

14

OpenStudy (anonymous):

hey @basilalter sorry i have some one like you at lies i thought he would gest and everything then the next day i dot a f sorry

OpenStudy (anonymous):

ok i have 7 more hich equation demonstrates the associative property? A. –(33 + 4) = –33 + (–4) B. (–16 + 9) + (–19) = –16 + (9 + (–19)) C. (–5 + 11) + 3 = (11 + (–5)) + 3 D. 2 + (–24) = –24 + 2

OpenStudy (anonymous):

im going with D

OpenStudy (anonymous):

@shadybubba it's all good.

OpenStudy (anonymous):

but idk, its been so long...

OpenStudy (anonymous):

ok thanks sorry so sorry

OpenStudy (anonymous):

ok well the next one is What value of z makes the statement true? –(z + (–17)) = 12 + 17

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

i do not know help me plz

OpenStudy (anonymous):

help

OpenStudy (anonymous):

eny one \